Two marbles (m = 0.050 kg each) moving on a table TOWARDS each other with the speed of 0.3 m/s collide and start moving away from each other.

a) What is the total momentum of this system before the collision?

V1 = __________ m/s, V2 = __________ m/s,

p1 = __________ kg m/s, p2 = __________ kg m/s,

ptotal = ____________ kg m/s

b) Determine the magnitude and direction of the velocity of each marble and the total linear momentum after the head-on elastic collision (hint: you may use the Collision lab to model this event. The initial POSITIONS of the balls don’t matter as long as they are some distance apart).

V1 = __________ m/s, V2 = __________ m/s,
Answer #1

a)
V1 = 0.3 m/s, V2 = -0.3 m/s

p1 = m*V1

= 0.05*0.3

= 0.015 kg.m/s

p2 = m*V2

= 0.05*(-0.3)

= -0.015 kg.m/s

p_total = p1 + p2

= 0.015 - 0.015

= 0 kg.m/s


b) as both marbles have equal mass their velocities are interchanged after the collision.

V1 = -0.3 m/s

V2 = 0.3 m/s